OpenStudy (ria23):

Can someone help me with this question? I keep getting a different answer. ;-; Simplify the expression below ln(e^3/x^4) A.)3e/4 B.) 3-x^4 C.) 3-4ln(x) D.) 3/x^4 I keep getting 3+ln(1/x^4)

OpenStudy (campbell_st):

well dividing within a log can be simpified to subtracting the logs \[\log(\frac{a}{b}) = \log(a) - \log(b)\] this is the 1st step needed for a solution. then simplify the individual logs if you can. you have correctly simplified ln(e^3) = 3 so its just the other term that needs to be corrected

OpenStudy (ria23):

3-4ln(x)?

OpenStudy (michele_laino):

@Ria23 that's right!

OpenStudy (ria23):

Awesome! thank yhu guys c:

OpenStudy (michele_laino):

@Ria23 thank you!